# Online Posterior Sampling with a Diffusion Prior

### v1

@@ 6 Experiments
Figure 2: Evaluation of $\tt DiffTS$ on three synthetic problems. The first row shows samples from the true (blue) and diffusion model (red) priors. The second row shows the regret of $\tt DiffTS$ and the baselines as a function of round $n$.
We conduct three experiments: synthetic problems in $2$ dimensions (Sections 6.2 and C.1), a recommender system (Section 6.3), and a classification problem (Section C.2). In addition, we conduct an ablation study in Section C.3, where we vary the number of training samples for the diffusion prior and the number of diffusion stages $T$.
@@ 6.1 Experimental Setup
We have four baselines. Three baselines are variants of contextual Thompson sampling [11, 3]: with an uninformative Gaussian prior ($\tt TS$), a learned Gaussian prior ($\tt TunedTS$), and a learned Gaussian mixture prior ($\tt MixTS$) [21]. The last baseline is diffusion posterior sampling ($\tt DPS$) of Chung et al. [12]. We implement all TS baselines as described in Section 5. The uninformative prior is $\mathcal{N}(\mathbf{0}_{d},I_{d})$. $\tt MixTS$ is used only in linear bandit experiments because the logistic regression variant does not exist. The TS baselines are chosen to cover various levels of prior information. Our implementation of $\tt DPS$ is described in Appendix D. We also experimented with frequentist baselines, such as $\tt LinUCB$ [1] and the $\varepsilon$-greedy policy. They performed worse than $\tt TS$ and thus we do not report them here.
Each experiment is set up as follows. First, the prior distribution of $\theta_{*}$ is specified: it can be synthetic or estimated from real-world data. Second, we learn this distribution from 10 0001000010\,00010 000 samples from it. In $\tt DiffTS$ and $\tt DPS$, we follow Appendix B. The number of stages is $T=100$ and the diffusion factor is $\alpha_{t}=0.97$. Since $0.97^{100}\approx 0.05$, most of the information in the training samples is diffused. The regressor in Appendix B is a $2$-layer neural network with ReLU activations. In $\tt TunedTS$, we fit the mean and covariance using maximum likelihood estimation. In $\tt MixTS$, we fit the Gaussian mixture using scikit-learn. All algorithms are evaluated on $\theta_{*}$ sampled from the true prior.
@@ 6.2 Synthetic Experiment
The first experiment is on three synthetic problems. Each problem is a linear bandit (Section 5) with $K=100$ actions in $d=2$ dimensions. The reward noise is $\sigma=2$. The feature vectors of actions are sampled uniformly at random from a unit ball. The prior distributions of $\theta_{*}$ are shown in Figure 2. The first is a mixture of two Gaussians and the last can be approximated well by a mixture of two Gaussians. $\tt MixTS$ is implemented with two mixture components, which can exactly represent the first prior and approximate the last one well.
Our results are reported in Figure 2. We observe two main trends. First, samples from the diffusion prior closely resemble those from the true prior. In such cases, $\tt DiffTS$ is expected to perform well and even outperforms $\tt MixTS$, because it has a better representation of the prior. We observe this in all problems. Second, $\tt DPS$ diverges as the number of rounds increases. This is because $\tt DPS$ relies on the score of the likelihood of observations (Section 7), which tends to infinity as the number of observations increases. This happens despite tuning (Appendix D). We report results on additional synthetic problems in Section C.1.
@@ 6.3 MovieLens Experiment
Figure 3: Evaluation of $\tt DiffTS$ on the MovieLens dataset.
In the second experiment, we learn to recommend an item to randomly arriving users. The problem is simulated using the MovieLens 1M dataset [27], with $1$ million ratings for 3 70637063\,7063 706 movies from 6 04060406\,0406 040 users. As a first step, we subtract the mean rating from all ratings and complete the sparse rating matrix $M$ by alternating least squares [14] with rank $d=5$. The learned factorization is $M=UV^{\top}$. The $i$-th row of $U$, denoted by $U_{i}$, represents user $i$. The $j$-th row of $V$, denoted by $V_{j}$, represents movie $j$. We use the item embeddings $V_{j}$ as unknown model parameters and the user embeddings $U_{i}$ as features of the actions.
We experiment with both linear and logistic bandits. In both, an item is initially chosen randomly from $V_{j}$ and $K=10$ random actions are chosen from $U_{i}$ in each round. In the linear bandit, the mean reward for recommending item $j$ to user $i$ is $U_{i}^{\top}V_{j}$. The reward noise is $\sigma=0.75$ and estimated from data. In the logistic bandit, the mean reward is $g(U_{i}^{\top}V_{j})$ where $g$ is a sigmoid.
Our MovieLens results are shown in Figure 3. We observe the same trends as in Section 6.2. First, samples from the diffusion prior closely resemble those from the true prior (Figure 3a). Since the problem is higher dimensional, we visualize the overlap using a UMAP projection [44]. Second, $\tt DiffTS$ has a significantly lower regret than all baselines, in both linear (Figure 3b) and logistic (Figure 3c) bandits. Finally, $\tt MixTS$ barely outperforms $\tt TunedTS$. We observe this trend consistently in higher dimensions, which motivated our work on online learning with more complex priors.
@@ 7 Related Work
There are two lines of related works: bandits with diffusion models and posterior sampling in diffusion models. We start with the former. Hsieh et al. [22] proposed Thompson sampling with a diffusion model prior for $K$-armed bandits. We consider a more general setting of contextual bandits and also develop new posterior sampling approximations for it. Aouali [4] proposed and analyzed contextual bandits with a linear diffusion model prior: $\mu_{t}(s_{t})$ in (5) is linear in $s_{t}$ and $q(s_{0})$ is a Gaussian. Due to this assumption, the resulting model is just a linear Gaussian model and not a general diffusion model, as in our work.
@@
The closest related work on posterior sampling in diffusion models is $\tt DPS$ of Chung et al. [12]. The key idea in $\tt DPS$ is to sample from a diffusion model posterior by adding the score of the likelihood of observations to it. In our notation, this is $\nabla\log p(h\mid\theta)$, where $p(h\mid\theta)$ is defined in Assumptions 1 and 2. Note that $\nabla\log p(h\mid\theta)$ grows linearly in $N$. Therefore, $\tt DPS$ becomes unstable as $N\to\infty$. See our empirical results in Section 6.2 and the implementation of $\tt DPS$ in Appendix D, which was tuned to improve its stability. Other methods for posterior sampling in diffusion models have been proposed recently: a sequential Monte Carlo approximation for the conditional reverse process [52], a variant of $\tt DPS$ with an uninformative prior [37], a pseudo-inverse approximation to the likelihood of evidence [47], and posterior sampling in latent diffusion models [40]. All of these methods rely on the score of the likelihood $\nabla\log p(h\mid\theta)$ and thus become unstable as the number of observations $N$ increases. Our posterior approximations do not have this issue because they are based on the product of prior and evidence distributions (Theorems 2 and 4), and thus gradient-free. They work well across different levels of uncertainty (Section 6) and do not require tuning.
@@
We also wanted to note that posterior sampling is a special from of guiding generation in diffusion models. Other approaches include conditional pre-training [15], a constraint in the reverse process [17], refining the null-space content [50], solving an optimization problem that pushes the reverse process towards evidence [46], and aligning the reverse process with the prompt [7].
@@ 8 Conclusions
We propose posterior sampling approximations for diffusion models priors. These approximations are contextual, and can be implemented efficiently in linear models and GLMs. We analyze them and evaluate them empirically on contextual bandit problems.
Our work focuses on developing novel posterior approximations rather than a regret analysis. This is because the existing approximations are unstable and may diverge in the online setting (Sections 7 and 6.2). A proper regret bound for $\tt DiffTS$ would require analyzing two errors. The first error arises because the reverse process reverses the forward process only approximately (Appendix B). The second error arises because our posteriors are only approximate (Section 4.3). That being said, prior works already provided general insights on the utility of complex priors. For instance, Russo and Van Roy [42] proved a $O(\sqrt{\Gamma H(A_{*})n})$ regret bound for a linear bandit, where $\Gamma$ is the maximum ratio of regret to information gain and $H(A_{*})$ is the entropy of the distribution of the optimal action under the prior. In short, the bound says that a lower entropy $H(A_{*})$, which reflects how concentrated the prior is, yields a lower regret.
Finally, we believe that our ideas can be extended beyond GLMs. The key idea in Section 4.4 is to use the Laplace approximation of the likelihood. This approximation can be computed exactly in GLMs. More generally though, it is a good approximation whenever the likelihood can be approximated well by a single Gaussian distribution. By the central limit theorem, under appropriate conditions, this is expected for any observation model when the number of observations is large.
Appendix A Proofs and Supplementary Lemmas
@@
A.1 Proof of Lemma 1
All derivations are based on basic rules of probability and the chain structure in Figure 1, and are exact. From Figure 1, the joint probability distribution conditioned on $H=h$ factors as
@@ A.1 Proof of Lemma 1
In the last step, we use that $p(h\mid s_{t})$ is a constant, since $s_{t}$ and $h$ are fixed, and that $p(h\mid s_{t-1},s_{t})=p(h\mid s_{t-1})$. Note that the last term $p(s_{t-1}\mid s_{t})$ is the conditional prior distribution. Let $t>1$. Then we rewrite the first term as
$p(h\mid s_{t-1})$ $=\int_{s_{0}}p(h,s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}$
$=\int_{s_{0}}p(h\mid s_{0},s_{t-1})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}$
@@
In the last equality, we use that our graphical model is a chain (Figure 1), and thus $p(h\mid s_{0},s_{t-1})=p(h\mid s_{0})$. Finally, we chain all identities and get that
@@
$p(h\mid s_{T})$ $=\int_{s_{0}}p(h,s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}$
$=\int_{s_{0}}p(h\mid s_{0},s_{T})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}$

### v2

@@ 6 Experiments
Figure 2: Evaluation of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ on three synthetic problems. The first row shows samples from the true (blue) and diffusion model (red) priors. The second row shows the regret of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ and the baselines as a function of round $n$.
We conduct three experiments: synthetic problems in $2$ dimensions (Sections˜6.2 and C.1), a recommender system (Section˜6.3), and a classification problem (Section˜C.2). In addition, we conduct an ablation study in Section˜C.3, where we vary the number of training samples for the diffusion prior and the number of diffusion stages $T$.
@@ 6.1 Experimental Setup
[…]
Each experiment is set up as follows. First, the prior distribution of $\theta_{*}$ is specified: it can be synthetic or estimated from real-world data. Second, we learn this distribution from 10 00010\,000 samples from it. In $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ and $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$, we follow Appendix˜B. The number of stages is $T=100$ and the diffusion factor is $\alpha_{t}=0.97$. Since $0.97^{100}\approx 0.05$, most of the information in the training samples is diffused. The regressor in Appendix˜B is a $2$-layer neural network with ReLU activations. In $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt TunedTS$, we fit the mean and covariance using maximum likelihood estimation. In $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt MixTS$, we fit the Gaussian mixture using scikit-learn. All algorithms are evaluated on $\theta_{*}$ sampled from the true prior. The regret is computed as defined in (9). All error bars are standard errors of the estimates.
@@ 6.2 Synthetic Experiment
The first experiment is on three synthetic problems. Each problem is a linear bandit (Section˜5) with $K=100$ actions in $d=2$ dimensions. The reward noise is $\sigma=1$. The feature vectors of actions are sampled uniformly at random from a unit ball. The prior distributions of $\theta_{*}$ are shown in Figure˜2. The first is a mixture of two Gaussians and the last can be approximated well by a mixture of two Gaussians. We implement $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt MixTS$ with two mixture components. Therefore, it can represent the first prior exactly and approximate the last one well.
Our results are reported in Figure˜2. We observe two main trends. First, samples from the diffusion prior closely resemble those from the true prior. In such cases, $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ is expected to perform well and even outperforms $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt MixTS$, because it has a better representation of the prior. We observe this in all problems. Second, $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ diverges as the number of rounds increases. This is because $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ uses an approximation based on the likelihood score (Section˜7), which is unstable when the score is high. This happens despite our best efforts to tune $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ (Appendix˜D). We report results on another three synthetic problems in Section˜C.1.
@@ 6.3 MovieLens Experiment
Figure 3: Evaluation of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ on the MovieLens dataset: (a) shows samples from the true (blue) and diffusion model (red) priors, (b) shows regret in the linear bandit, and (c) shows regret in the logistic bandit.
In the second experiment, we learn to recommend an item to randomly arriving users. The problem is simulated using the MovieLens 1M dataset [28], with one million ratings for 3 7063\,706 movies from 6 0406\,040 users. We subtract the mean rating from all ratings and complete the sparse rating matrix $M$ by alternating least squares [14] with rank $d=5$. The learned factorization is $M=UV^{\top}$. The $i$-th row of $U$, denoted by $U_{i}$, represents user $i$. The $j$-th row of $V$, denoted by $V_{j}$, represents movie $j$. We use movie embeddings $V_{j}$ as model parameters and user embeddings $U_{i}$ as features of the actions. The movies are items.
We experiment with both linear and logistic bandits. In both, an item is initially chosen randomly from $V_{j}$ and $K=10$ actions are chosen randomly from $U_{i}$ in each round. In the linear bandit, the mean reward of item $j$ for user $i$ is $U_{i}^{\top}V_{j}$. The reward noise is $\sigma=0.75$, and we estimate it from data. In the logistic bandit, the mean reward is $g(U_{i}^{\top}V_{j})$, where $g$ is a sigmoid.
Our MovieLens results are reported in Figure˜3 and we observe similar trends to Section˜6.2. First, samples from the diffusion prior closely resemble those from the true prior (Figure˜3a). Since the problem is higher dimensional, we visualize the overlap using UMAP [45]. Second, $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ has a lower regret than all baselines, in both linear (Figure˜3b) and logistic (Figure˜3c) bandits. Finally, $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt MixTS$ barely outperforms $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt TunedTS$. We observe this trend consistently in higher dimensions, and this motivated our work on online learning with more complex priors.
@@ 7 Related Work
We start with reviewing related works on bandits with diffusion models. Hsieh et al. [23] proposed Thompson sampling with a diffusion model prior for $K$-armed bandits. There are multiple technical differences from our work. First, the diffusion model in Hsieh et al. [23] is over scalars representing individual arms. Our model is over vectors representing model parameters, and thus can be applied to contextual bandits. Second, the approximations are different. In stage $t$, Hsieh et al. [23] sample from two distributions: the conditional prior and the distribution of the diffused empirical mean up to stage $t$. Then they take a weighted sum of the samples. We sample only once, from the posterior distribution that combines the conditional prior in stage $t$ and likelihood. Therefore, the method of Hsieh et al. [23] can be viewed as a non-contextual variant of our method, where posterior sampling is done by weighting samples from the prior and empirical distributions. Finally, Hsieh et al. [23] do not analyze their approximation.
Aouali [5] proposed and analyzed contextual bandits with a linear diffusion model prior: $\mu_{t}(s_{t})$ in (5) is linear in $s_{t}$ and $q(s_{0})$ is a Gaussian. Therefore, this model is a linear Gaussian model and not a general diffusion model, as in our work.
The closest related work on posterior sampling in diffusion models is $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ of Chung et al. [12]. The key idea in $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ is to sample from the posterior distribution using the likelihood score $\nabla\log p(h\mid\theta)$, where $p(h\mid\theta)$ is the likelihood (Assumptions˜1 and 2). Note that $\nabla\log p(h\mid\theta)$ grows linearly in $N$ because the history $h$ in $p(h\mid\theta)$ involves $N$ terms. Therefore, $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ becomes unstable as $N\to\infty$. We show it empirically in Section˜6.2 and discuss the implementation of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ in Appendix˜D, which was tuned to improve its stability.
@@
Many other posterior sampling methods for diffusion models have been proposed recently: a sequential Monte Carlo approximation for the conditional reverse process [53], a variant of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ with an uninformative prior [38], a pseudo-inverse approximation to the likelihood of evidence [48], and posterior sampling in latent diffusion models [41]. All of these methods rely on the likelihood score $\nabla\log p(h\mid\theta)$ and thus become unstable as the number of observations $N$ increases. Our posterior approximations do not have this issue because they are based on the product of prior and evidence distributions (Theorems˜2 and 4), and thus gradient-free. They work well across different levels of uncertainty (Section˜6) and do not require tuning.
@@
We note that posterior sampling is a special form of inference-time guidance in diffusion models. Other approaches are conditional pre-training [15], a constraint in the reverse process [18], refining the null-space content [51], solving an optimization problem that pushes the reverse process towards evidence [47], and aligning the reverse process with the prompt [7].
@@ 8 Conclusions
We propose posterior sampling approximations for diffusion models priors. These approximations are contextual, and can be implemented efficiently in linear models and GLMs. We analyze them and evaluate them empirically on contextual bandit problems. Our method has two main limitations.
Computational cost. The cost of posterior sampling in $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$ with $T$ stages is about $T$ times higher than that of posterior sampling with a Gaussian prior (Section˜2). We validate it empirically in Section˜6.2. We plot the sampling time as a function of $T$ in Figure˜6c (Section˜C.3).
Learning cost and hyper-parameter tuning. In all experiments, the number of diffusion stages is $T=100$ and the diffusion rate is set such that most of the signal diffuses. The regressor is a $2$-layer neural network and we learn it from 10 00010\,000 samples from the prior. These settings resulted in stable performance in all our experiments (Section˜6). However, they clearly impact the performance. We plot the regret as a function of the number of training samples in Figure˜6a and as a function of $T$ in Figure˜6b. When $T$ or the number of training samples is small, $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ performs very similarly to posterior sampling with a Gaussian prior. In summary, there is no benefit in these cases.
Future work. We develop novel posterior approximations rather than bounding their regret. This is because the existing approximations are unstable and may diverge in the online setting (Sections˜7 and 6.2). We believe that a proper regret analysis of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ is possible and would require bounding two errors. The first error arises because the reverse process does not reverse the forward process exactly (Appendix˜B). The second error arises because our posterior distributions are approximate (Section˜4.3). One possibility is to start with prior works that already showed the utility of complex priors. For instance, Russo and Van Roy [44] proved a $O(\sqrt{\Gamma H(A_{*})n})$ regret bound for a linear bandit, where $\Gamma$ is the maximum ratio of regret to information gain and $H(A_{*})$ is the entropy of the distribution of the optimal action under the prior. This bound holds for any prior and says that a lower entropy $H(A_{*})$, which corresponds to more informative priors, yields a lower regret.
We also believe that our ideas can be extended beyond GLMs. The key idea in Section˜4.4 is to use the Laplace approximation of the likelihood. This approximation can be computed exactly in GLMs. More generally though, it is a good approximation whenever the likelihood can be approximated well by a single Gaussian distribution. By the central limit theorem, under appropriate assumptions, this is expected for any observation model when the number of observations is large.
Appendix A Proofs and Supporting Lemmas
@@
A.1 Proof of Lemma˜1
All derivations are based on basic rules of probability and the chain structure in Figure˜1, and are exact. From Figure˜1, the joint probability distribution conditioned on $H=h$ factors as
@@ A.1 Proof of Lemma˜1
In the last step, we use that $p(h\mid s_{t})$ is a constant, since $s_{t}$ and $h$ are fixed, and that $p(h\mid s_{t-1},s_{t})=p(h\mid s_{t-1})$. Note that the last term $p(s_{t-1}\mid s_{t})$ is the conditional prior distribution. When $t>1$, the first term can be expressed as
$p(h\mid s_{t-1})$ $=\int_{s_{0}}p(h,s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}=\int_{s_{0}}p(h\mid s_{0},s_{t-1})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}$
@@
In the last equality, we use that our graphical model is a chain (Figure˜1), and thus $p(h\mid s_{0},s_{t-1})=p(h\mid s_{0})$. Finally, we chain all identities and get that
@@
$p(h\mid s_{T})$ $=\int_{s_{0}}p(h,s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}=\int_{s_{0}}p(h\mid s_{0},s_{T})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}$